# Dissipative qutrit-mediated stable charging

### v1

@@ Appendix A The effective master equation
$\dot{\rho}(t)=$ $-i[(\tilde{H}+V)\rho(t)-\rho(t)(\tilde{H}^{\dagger}+V)]$ (30)
@@
In the weak-driving regime Reiter and Sørensen (2012), the subspace spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ with $n<N$ in Fig. 2(a) can be divided into the rapidly-decaying exited subspace spanned by $\{|hn\rangle,|e(n+1)\rangle\}$ and the ground subspace spanned by $\{|gn\rangle,|g(n+1)\rangle\}$. With respect to Fig. 2(a), the non-Hermitian Hamiltonian defined in Eq. (28) is given by
@@
involving only with the excited subspaces. The weak external driving $V$ leads to the excitation process from the ground subspaces to the excited subspaces with a Hermitian Hamiltonian
$V_{n}=V_{+,n}+V^{\dagger}_{+,n}=\Omega|hn\rangle\langle gn|+{\rm H.c.}.$ (32)
@@
$\dot{\tilde{\rho}}(t)=$ $-i[\tilde{V}(t)\tilde{\rho}(t)-\tilde{\rho}(t)\tilde{V}^{\dagger}(t)]+\tilde{L}_{hg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{hg}(t)$ (34)
@@
where $\tilde{\rho}(t)=\tilde{U}(t)\rho(t)\tilde{U}^{\dagger}(t)$ is the density matrix in the rotating frame. The other operators are transformed accordingly with $\tilde{\mathcal{O}}(t)=\tilde{U}(t)\mathcal{O}\tilde{U}^{-1}(t)$, $\mathcal{O}=V$, $L_{hg}$, $L_{eg}$, and $L_{he}$.
With respect to the main process in Fig. 2(a), the transformed master equation (34) in the subspace can be written as
@@
$+\tilde{L}_{hg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{hg,n}(t)$
$+\tilde{L}_{eg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{eg,n}(t),$
@@
$\tilde{V}_{n}(t)$ $=\Omega(e^{i\tilde{H}_{n}t}|hn\rangle\langle gn|+|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t}),$ (36)
@@
Then we vectorize the density matrix of each subspace to be a $4^{2}\times 1$ vector denoted by $|\tilde{\rho}_{n}(t)\rangle\rangle$. The equation of motion in this representation is given by
@@
Her, $\mathbf{L}(t)$ is the $4^{2}\times 4^{2}$ superoperator in the form of
@@
$\mathcal{V}_{n}(t)=-i(\mathbb{I}\otimes\tilde{V}_{n}(t)-\tilde{V}_{n}^{*}(t)\otimes\mathbb{I}),$
@@
where $\mathbb{I}$ is the $4\times 4$ identity operator in the subspace Campaioli et al. (2024b). $\mathcal{V}(t)$ and $\mathcal{J}(t)$ represent the non-Hermitian evolution part and the quantum jump part, respectively. The formal solution of Eq. (39) is
@@
In the large detuning condition, i.e., $\Delta+\delta\gg\Omega$, the transformation operator $U(t)=e^{i\tilde{H}_{n}(t)}\propto\exp[i(\tilde{\Delta}+\tilde{\delta})t/2]$ is highly oscillating with time. By iteratively substituting Eq. (41) into Eq. (39), neglecting the fast oscillating terms $\mathcal{V}(t)|\tilde{\rho}_{n}(0)\rangle\rangle$, and taking the Markovian approximation Shao et al. (2017), the equation of motion can written into a Dyson series:
@@
$=$ $\left[\mathbf{L}^{(2)}_{{\rm eff},n}(t)+\mathbf{L}^{(3)}_{{\rm eff},n}(t)+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$
@@
$+\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\bigg{]}\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle.$
up to the second order in driving intensity $\Omega$. By substituting Eq. (36) into the first term of the Eq. (43), we can obtain the effective superoperator for a non-Hermitian evolution:
[…]
$=$ $-\left[P_{g}\otimes V_{+,n}^{\dagger}(i\tilde{H}_{n})^{-1}V_{+,n}+V_{+,n}^{\dagger}(-i\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes P_{g}\right],$
[…]
@@
By substituting Eqs. (36)-(38) into the second term of Eq. (43), we can obtain the effective superoperator for quantum jumps:
[…]
@@
$L_{hg,n}^{\text{eff}}=$ $L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|,$ (47)
$L_{eg,n}^{\text{eff}}=$ $L_{eg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|.$ (48)
Therefore according to Eq. (43), the equation of motion in the ground state subspaces reads
$P_{g}\dot{\rho}_{n}(t)P_{g}=-i\left[\mathcal{H}^{\rm eff}_{n}\rho_{n}(t)-\rho_{n}(t)(\mathcal{H}^{\rm eff}_{n})^{\dagger}\right]$ (49)
@@
$P_{g}\dot{\rho}_{n}(t)P_{g}=$ $-i\left[H^{\text{eff}}_{n},\rho_{n}(t)\right]+\mathcal{L}[L_{hg,n}^{\rm eff}]\rho_{n}(t)$ (50)
@@
$h_{n}^{\rm eff}=\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}\sum_{k=hg,eg}(L^{\rm eff}_{k,n})^{\dagger}L^{\rm eff}_{k,n}$ (52)
$=$ $\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}V_{+,n}^{\dagger}(\tilde{H}_{n}^{\dagger})^{-1}\sum_{k=hg,eg,he}L_{k,n}^{\dagger}L_{k,n}\tilde{H}_{n}^{-1}V_{+,n}$
@@
$=$ $H^{\rm eff}_{n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n},$
where $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The non-Hermitian part represents the leakage from $|gn\rangle$ to $|en\rangle$ with the jump operator $L^{\rm eff}_{he,n}=L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$. It is negligible comparing to the spontaneous decay $\gamma_{eg}$ under the condition $\Delta,\gamma_{hg}\gg\gamma_{he},\Omega$.
For the uppermost ground-state subspace spanned by $\{|gN\rangle,|hN\rangle\}$, the non-Hermitian Hamiltonian reads
@@
$L^{\rm eff}_{hg,N}=$ $L_{hg,N}\tilde{H}_{N}V_{+,N}=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}|gN\rangle\langle gN|,$ (54)
@@
The leading-order correction induced by the decay $\gamma_{he}$ is described in Fig. 2(b). The process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|gn\rangle$ gives rise to a pure dephasing effect on $|gn\rangle$, which is associated with a higher order term $\mathbf{L}^{(4)}_{{\rm eff},n}$ in Eq. (42)
[…]
×$\times\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g},$
@@
with $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The dephasing rate turns out to be
@@
×$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle$
using the identity $e^{-i\tilde{H}^{*}t}\otimes e^{i\tilde{H}t}=\exp(-i\tilde{H}^{*}t\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}t)$. Similarly, the process $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ in Fig. 2(b) gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$, the relevant jump superoperator is given by
@@
×$\times\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g}.$
@@
×$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle.$
Under the parametric settings that $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$, it is important to observe that the dephasing rate given by Eq. (58) is negligible in comparison to that in Eq. (47). Similarly, the effective transition rate from $|gn\rangle$ to $|g(n-1)\rangle$ given by Eq. (61) is negligible in comparison to that from $|gn\rangle$ to $|g(n+1)\rangle$ in Eq. (48) under the condition $\gamma_{eg}\approx\gamma_{he}\ll\gamma_{hg}$ and $\Delta\gg\delta$ due to our optimization condition $g^{2}A_{n}^{2}\approx|\tilde{\Delta}\tilde{\delta}|$.

### v2

@@ Appendix A Effective master equation
$\dot{\rho}(t)=$ $-i\left[(\tilde{H}+V)\rho(t)-\rho(t)(\tilde{H}^{\dagger}+V)\right]$ (30)
@@
In the weak-driving regime Reiter and Sørensen (2012), the subspace spanned by $\{|gn\rangle,|hn\rangle,|e(n+1)\rangle,|g(n+1)\rangle\}$ with $n<N$ in Fig. 2(a) can be divided into the rapidly decaying exited subspace spanned by $\{|hn\rangle,|e(n+1)\rangle\}$ and the ground subspace spanned by $\{|gn\rangle,|g(n+1)\rangle\}$. With respect to Fig. 2(a), the non-Hermitian Hamiltonian defined in Eq. (28) is given by
@@
involving only with the excited subspaces. The weak external driving $V$ connects the ground subspaces to the excited subspaces with a Hermitian Hamiltonian
$V_{n}=V_{+,n}+V^{\dagger}_{+,n}=\Omega|hn\rangle\langle gn|+{\rm H.c.}$ (32)
@@
$\dot{\tilde{\rho}}(t)=$ $-i\left[\tilde{V}(t)\tilde{\rho}(t)-\tilde{\rho}(t)\tilde{V}^{\dagger}(t)\right]+\tilde{L}_{hg}(t)\tilde{\rho}(t)\tilde{L}^{\dagger}_{hg}(t)$ (34)
@@
where $\tilde{\rho}(t)=\tilde{U}(t)\rho(t)\tilde{U}^{\dagger}(t)$ is the density matrix in the rotating frame. The other operators are transformed accordingly to be $\tilde{\mathcal{O}}(t)=\tilde{U}(t)\mathcal{O}\tilde{U}^{-1}(t)$, $\mathcal{O}=V$, $L_{hg}$, $L_{eg}$, and $L_{he}$.
With respect to the main transitions in Fig. 2(a), the transformed master equation (34) in the $n$th subspace can be written as
@@
$+$ $\tilde{L}_{hg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{hg,n}(t)+\tilde{L}_{eg,n}(t)\tilde{\rho}_{n}(t)\tilde{L}^{\dagger}_{eg,n}(t),$
@@
$\tilde{V}_{n}(t)$ $=\Omega\left(e^{i\tilde{H}_{n}t}|hn\rangle\langle gn|+|gn\rangle\langle hn|e^{-i\tilde{H}_{n}t}\right),$ (36)
@@
Then we vectorize the density matrix of each subspace to be a $4^{2}\times 1$ vector denoted by $|\tilde{\rho}_{n}(t)\rangle\rangle$. The equation of motion in the Hilbert-Schmidt space is given by
@@
Here $\mathbf{L}(t)$ is the $4^{2}\times 4^{2}$ superoperator in the form of
@@
$\mathcal{V}_{n}(t)=-i\left[\mathbb{I}\otimes\tilde{V}_{n}(t)-\tilde{V}_{n}^{*}(t)\otimes\mathbb{I}\right],$
@@
where $\mathbb{I}$ is the $4\times 4$ identity operator in the subspace Campaioli et al. (2024b). $\mathcal{V}(t)$ and $\mathcal{J}(t)$ represent the non-unitary evolution part and the quantum jump part, respectively. The formal solution of Eq. (39) is
@@
Under the large detuning condition, i.e., $\Delta+\delta\gg\Omega$, the transformation operator $\tilde{U}_{n}(t)=e^{i\tilde{H}_{n}t}\propto\exp[i(\tilde{\Delta}+\tilde{\delta})t/2]$ is highly oscillating with time. By iteratively substituting Eq. (41) into Eq. (39), neglecting the fast oscillating terms $\mathcal{V}(t)|\tilde{\rho}_{n}(0)\rangle\rangle$, and taking the Markovian approximation Shao et al. (2017), the equation of motion can written into a perturbative Dyson series:
@@
$=$ $\left[\mathbf{L}^{(2)}_{{\rm eff},n}(t)+\mathbf{L}^{(3)}_{{\rm eff},n}(t)+\mathbf{L}^{(4)}_{{\rm eff},n}+\cdots\right]|\tilde{\rho}_{n}(t)\rangle\rangle.$
@@
$+\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\bigg{]}\mathcal{P}_{g}|\tilde{\rho}_{n}(t)\rangle\rangle$
up to the second order in the driving intensity $\Omega$. By substituting Eq. (36) into the first term of the Eq. (43), we can obtain the effective superoperator for a nonunitary evolution:
$\mathcal{P}_{g}\mathcal{V}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\mathcal{P}_{g}$ (44)
$=$ $-P_{g}\otimes V_{+,n}^{\dagger}(i\tilde{H}_{n})^{-1}V_{+,n}$
$-V_{+,n}^{\dagger}(-i\tilde{H}_{n}^{*})^{-1}V_{+,n}\otimes P_{g},$
where we used the identity $\int dte^{i\tilde{H}t}=(i\tilde{H})^{-1}e^{i\tilde{H}t}$. In the original representation for density matrix, it corresponds to the non-Hermitian effective Hamiltonian
@@
By substituting Eqs. (36)-(38) into the second term of Eq. (43), one can obtain the effective superoperators for quantum jump:
$\mathcal{P}_{g}\mathcal{J}_{n}(t)\int_{0}^{t}dt_{1}\mathcal{V}_{n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\mathcal{P}_{g}$ (46)
@@
$L_{hg,n}^{\rm eff}=$ $L_{hg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{hg}}\Omega\tilde{\delta}}{\tilde{\Delta}\tilde{\delta}-g^{2}A_{n}^{2}}|gn\rangle\langle gn|,$ (47)
$L_{eg,n}^{\rm eff}=$ $L_{eg,n}(\tilde{H}_{n})^{-1}V_{+,n}=\frac{\sqrt{\gamma_{eg}}\Omega gA_{n}}{g^{2}A_{n}^{2}-\tilde{\Delta}\tilde{\delta}}|g(n+1)\rangle\langle gn|.$ (48)
Thus according to Eq. (43), the equation of motion in the ground-state subspaces reads
$P_{g}\dot{\rho}_{n}(t)P_{g}\approx-i\left[\mathcal{H}^{\rm eff}_{n}\rho_{n}(t)-\rho_{n}(t)(\mathcal{H}^{\rm eff}_{n})^{\dagger}\right]$ (49)
@@
$P_{g}\dot{\rho}_{n}(t)P_{g}\approx$ $-i\left[H^{\rm eff}_{n},\rho_{n}(t)\right]+\mathcal{L}[L_{hg,n}^{\rm eff}]\rho_{n}(t)$ (50)
@@
$\mathcal{H}^{\rm eff}_{n}+\frac{i}{2}\sum_{k=hg,eg}\left(L^{\rm eff}_{k,n}\right)^{\dagger}L^{\rm eff}_{k,n}$ (52)
@@
$=$ $H^{\rm eff}_{n}-\frac{i}{2}(L^{\rm eff}_{he,n})^{\dagger}L^{\rm eff}_{he,n}.$
The non-Hermitian part represents the leakage from $|gn\rangle$ to $|en\rangle$ with the jump operator $L^{\rm eff}_{he,n}=L_{he,n}\tilde{H}_{n}^{-1}V_{+,n}$. It is negligible comparing to the spontaneous decay $\gamma_{eg}$ under the condition $\Delta,\gamma_{hg}\gg\gamma_{he},\Omega$.
For the uppermost subspace spanned by $\{|gN\rangle,|hN\rangle\}$, the non-Hermitian Hamiltonian reads
@@
$L^{\rm eff}_{hg,N}=$ $L_{hg,N}(\tilde{H}_{N})^{-1}V_{+,N}=\frac{\sqrt{\gamma_{hg}}\Omega}{\tilde{\Delta}}|gN\rangle\langle gN|,$ (54)
@@
The leading-order discharging induced by the decay $\gamma_{he}$ is described in Fig. 2(b). The three-step transition $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|gn\rangle$ gives rise to a pure dephasing effect on $|gn\rangle$. It is associated with a higher order term $\mathbf{L}^{(4)}_{{\rm eff},n}$ in Eq. (42)
$J^{\text{eff}*}_{eg,n}\otimes J^{\rm eff}_{eg,n}=$ $\mathcal{P}_{g}\tilde{L}^{*}_{eg,n-1}(t)\otimes\tilde{L}_{eg,n-1}(t)\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})$ (56)
$\times\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g},$
@@
with $L_{he,n}=\sqrt{\gamma_{he}}|en\rangle\langle hn|$. The dephasing rate turns out to satisfy
@@
$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle,$
using the identity $e^{-i\tilde{H}^{*}t}\otimes e^{i\tilde{H}t}=\exp(-i\tilde{H}^{*}t\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}t)$. Similarly, the four-step transition $|gn\rangle\rightarrow|hn\rangle\rightarrow|en\rangle\rightarrow|h(n-1)\rangle\rightarrow|g(n-1)\rangle$ in Fig. 2(b) gives rise to an effective transition from $|gn\rangle$ to $|g(n-1)\rangle$, the relevant jump superoperator is given by
@@
$\times\int_{0}^{t}dt_{1}\mathcal{J}_{he,n}(t_{1})\int_{0}^{t_{1}}dt_{2}\mathcal{V}_{n}(t_{2})\int_{0}^{t_{2}}dt_{3}\mathcal{V}_{n}(t_{3})\mathcal{P}_{g}.$
@@
$\times\left(-i\tilde{H}^{*}_{n-1}\otimes\mathbb{I}+\mathbb{I}\otimes i\tilde{H}_{n-1}\right)^{-1}|en\rangle\otimes|en\rangle.$
[…]